# Tree networks of real-world data: analysis of efficiency and spatiotemporal scales

### v4

@@ 3 Methods
Figure 3: Search method based on relative distance. In (a) the growth process is illustrated. At each step, a new pattern is linked to either the current parent pattern, $P$, or to one of its children, $C_{i}$. The parameter $r$ modulates the region of attraction of the child, indicated by the gray circle around $C_{1}$, and the one of the parent node. Points falling within the distance $r\cdot d(P,C_{i})$ are linked to the child, otherwise, an edge will be defined between the parent node $P$ and the new pattern. In (b) the search method based on relative distance is illustrated. Once a search pattern arrives at node $P$ it must choose to search either the subtree of $C_{1}$ or $C_{2}$. The choice is based on minimizing the distance to the expected patterns contained in the subtree of $C_{i}$. In the depicted case, patterns closer to the search pattern $S$ are expected to be in the subtree of $C_{2}$. Adapted from [33].
3.1 Pattern-matching efficiency
@@
After the tree formation described in Section 1, the Portegys algorithm enters the search phase, where the best match to an unknown search pattern, $S$, is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list while providing a considerable speed up due to the hierarchical organization into a tree-based structure. The growth process, described in Section 1 and depicted in panel (a) of Figure 3, ensures that subtrees of a given node have on average decreasing distances for decreasing hierarchical levels. This provides the possibility of using a search technique based on relative distances. The search method is depicted in panel (b) of Figure 3 and is an iterative process that begins at the root-node pattern. Once a search pattern, $S$, arrives at a certain parent node, $P$, it must decide to continue to search for the best match into one of its children subtree, $C^{\prime}_{1}$, $C^{\prime}_{2}$ in Figure 3. The choice is taken minimizing the Portegys search-distance: $d(S,C_{i})_{search}=d(S,C_{1})-r\cdot d(P,C_{i})=d(S,C^{\prime}_{i})$, which quantifies the distance between the $S$-pattern and the potential patterns appearing in the subtree of each child $C_{i}$, collectively indicated by the gray circle $C^{\prime}_{i}$ in panel (b). Where $d(\cdot)$ indicates any metric distance equipped with semi-positivity, symmetry, and triangular inequality properties. If $d(S,C_{i})_{search}$ is less than zero then it is set to zero. In panel (b) the bold segments in red indicate the search distance to each child node and equivalently the expected distance between the unknown pattern $S$ and their expected neighborhoods of patterns $C^{\prime}_{i}$.
@@
It is natural to define a pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
@@
the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$, where $N$ is the number of nodes. For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$, for lower values of $r$ the number of computed distances decreases, potentially decreasing the class-match accuracy. When defining the pattern-matching efficiency we are making the simplified assumption that the pattern-matching algorithm runs on a physical substrate such that visiting any node has a fixed cost in terms of energy or terms of general resources. Hence, the efficiency $\theta$ quantifies the trade-off between the cost of used resources and the gain in performance, as depicted in panel (c, f) of Figure 4 for the MNIST and CIFAR10 respectively. Note in panel (d) that the CIFAR10 dataset has a peak for accuracy which we interpret by saying that for complex datasets (see also the case of the FashionMNIST in Section S7) some network structures do possess the ability to better encode information of the "world" in their traits. We postpone the discussion of this finding to Sections 5 and 6.
@@
Figure 4: Pattern-matching efficiency in (c, f) measured over the NIST and CIFAR10 respectively. The efficiency $\theta$ is defined as the trade-off between the classification accuracy, (a, d), and the logarithm of the number of distances computed to find the best match, (b, e). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency $\theta$. The shaded grey area designates the standard deviation. The purple shaded area in panel (c, f) indicates a tolerance $r$-domain of optimal efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by the yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST and CIFAR10 datasets. The 10-digit classes are equally represented both in the train- and test sets.
4 Phase portrait
Figure 5: Multiple phase transitions in hierarchical tree networks. In (a) the specific heat, $C$, across the range $r\in[0,2]$. At $r\sim 1.35$, indicated by the dashed white line, the hierarchical network transits from the star graph to the amorphous configuration with the emergence of scale-invariant traits evidenced by a plateau in the specific heat. The scale-invariant region shows a second phase transition from amorphous to regular 1d-lattice at $r\simeq 1.05$ indicated by the dotted white line. For $r>0.35$, the network is always scale-invariant with decreasing spectral dimension up to the transition point, above which the spectral dimension saturates to $d_{s}=1=2\cdot C_{0}$, where $C_{0}$ is the plateau height in (c). Note that at $r\sim 0.8$ the spectral dimension is $d_{s}=4/3$ as expected for random trees with finite variance in the node degree. In panel (c) the entropy (blue) and the specific heat (red) for $r=\{0.35,~{}0.40,~{}0.70,~{}0.80,~{}0.90,2.00\}$. Vertical blue and green lines indicate the $\tau$-scale of the first and the second peak, $\tau^{(1,2)}_{peak}$, of the specific heat respectively. In (d), the $\tau$-scales of the two specific heat peaks are illustrated with blue and green dots respectively delimiting the three dynamical phases: the segregated phase (grey), the integrated phase (light yellow) and the green region where self-similar structures are integrated at a constant-rate. In (b) exemplary hierarchical trees of $N=20$ for growing $r$ from left to right. All curves in this figure are realized for networks of $N=500$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations.
Figure 6: Order parameters. In (a), the modularity, $Q$, resolves the transition involving the mesoscale structure. In (b), the normalized diameter, $\hat{d}=d/m$ resolves the large-scale transition. Their respective variances (dashed lines) identify the transition points. In (c), the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks quantifies the normalized (log)number of informationally connected nodes across the scale-invariant $\tau$-domain, i.e. the number of nodes included in the scale-invariant region. Note the discontinuity at the transition from star to amorphous. Curves in (a-b) are realized for networks of $N=500,~{}1000$ randomly sampled from the FashionMNIST dataset and averaged over 50 independent realizations if not specified elsewhere. Curves in (c) are for networks of $N=500$.
@@ 4 Phase portrait
The second change of phase is the transition from the star graph to the amorphous configuration at $r\sim 0.35$. Here, the emergence of scale-invariant structural properties is reflected by the progressively larger specific heat plateau in between the two $C$-peaks, as shown in (a) and (c) of Figure 5. For $r\rightarrow 2$, the plateau would include the full 1d-lattice in the thermodynamic limit $N\rightarrow\infty$.
@@
Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. After the phase transition, a second peak of the specific heat emerges at larger $\tau$ scales. This second peak of $C$ in Figure 5.(c) identifies the scale at which the information is integrated between the distinct modules. In other words, the second peak reflects the "degree" of the large-scale modules, in much analogy to the first peak, which reflects the presence of hubs at the local connectivity scale of the nodes. Note that we exclude both the clusters, i.e. triangles, and higher-order cycles from our discussion as the tree networks we address are purely loop-less.
@@
This second change of phase is the one mainly investigated in this work as the peaks in the thermodynamic and pattern-matching efficiencies always occur in the scale-invariant region of the phase plane, and in some cases, they occur in the vicinity of this transition border $r\sim 0.35$. This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network. Indeed, we observe in Figure 6.(b) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition. Nevertheless, a phase transition exists and an appropriate order parameter needs to be employed to identify it.
@@
This phase change occurs at smaller resolution scales with respect to the first phase transition we discussed. Specifically, it involves the mesoscales properties changing from the star graph to the scale-invariant regime (see the bifurcation in the specific heat in Figure 5.(a),(c)). Thus, we can identify this transition using as an order parameter the modularity [35], $Q$, of the tree which we show in panel (a) of Figure 6. We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in [36]. We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor. This motivates the choice of a mesoscale $\tau\sim 10$ employed in Section 5 that can resolve structural changes from the star graph phase to the amorphous scale-invariant phase.
@@
As a second possible order parameter, for the mesoscale phase-transition, we show in Figure 6.(c) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ peaks. This order parameter counts the normalized (log)number of informationally connected nodes across the $\tau$-scales of the specific heat plateau. In simple terms, we count the fraction of nodes included in the scale-invariant region. We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
@@
As already discussed in Section 2.1, also along the $\tau$-direction a phase transition exists [22, 3] delimited by the two peaks of $C$ in $\tau^{(1,2)}_{peak}$ and, in between the two, we can distinguish three dynamical phases (see Figure 5.(d)). For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase. Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [3, 23].
5 Results
Figure 7 shows the pattern-matching efficiency defined in Equation 6 and the thermodynamic efficiency of Equation 5 for the four investigated datasets.
Panels (a)-(d) depict the thermodynamic efficiency, $\eta$, and panels (e)-(h) depict the pattern-matching efficiency, $\theta$. The vertical dashed lines indicate the peaks of the efficiencies, while the shaded purple regions in (e)-(h) indicate a $r$-domain tolerance for $\theta$. The tolerance domain is defined as those $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the best trade-off.
Note the consistency in the global maxima of $\theta$ and $\eta$ in terms of optimal $r$ for the NIST and the FashionMNIST. See the $r$-range spanned by the purple-shaded areas in panels (e)-(h) and the vertical dashed lines in panels (a)-(d). This consistency in the maxima suggests that the two efficiencies might be related, for the specific context of this algorithm. In the Section 6 we discuss this matter.
The topological complexity of the networks increases at a critical $r$ from the trivial traits of a star graph to the fractal structure of the amorphous network. Thus, the emergence of structure is reflected by the rise of a second peak in $C$ together with a roughly $C=C_{0}>0$ plateau in between the first and the last peak as is illustrated in panels (i)-(l) for the investigated dataset. Such transition is also depicted in Figure 8 by the bifurcation of the blue and green curves for $\tau^{(1)}_{peak}$ and $\tau^{(2)}_{peak}$ indicating the lower and upper branches in panels (i)-(l) of Figure 7. See also Section S4 which illustrates the spectral entropy and specific heat curves for the four datasets independently depicted for each $r$-value.
We note that the specific heat of the MNIST has a third large peak in between $\tau^{(1)}$ and $\tau^{(2)}$ (see panel (k) in Figure 7). This is an intrinsic characteristic of the dataset, as it is only mitigated for larger networks, as shown in Section S2 of the Supplementary. This intermediate peak is sufficiently strong to be reflected on a local maximum of $\eta$, in Figure 7.(c). Although this local maximum in $\eta$ is not detected by $\theta$ in Figure 7(g), it is reflected in its standard deviation, allowing us to include the $r$-range indicated by the shaded purple area when we compare $\eta$ and $\theta$ peaks in terms of $r$. This shaded region corresponds to the $\eta$-peak-to-peak $r$-range in Figure 7.(c).
Remarkably, the thermodynamic efficiency is computed at the specific resolution scale $\tau\sim 10$ involved in the mesoscale structural transition from star graph to amorphous scale-invariant discussed in Sections 4 and 6. Such critical resolution scale is identified by the bifurcation of the specific heat peaks in panels (i)-(l) of Figure 7 and in panels (a)-(d) of Figure 8. In the seminal work [1], it is suggested to compute the thermodynamic efficiency, $\eta$, using as a reference the $\tau$-scale equal to the inverse of the Fiedler eigenvalue, i.e. the smallest non-zero Laplacian eigenvalue $\lambda_{2}=1/\tau_{diff}$, and logarithmically divide $\tau_{diff}$ to define the large propagation scales, the middle scales, and the small scales. In fact, this eigenvalue should provide the time scale where the diffusion process has reached equilibrium.
Nevertheless, on the one hand, $\tau_{diff}$ only approximates the end of the diffusion process which is better reflected by $\tau^{(2)}_{peak}$. Panels (a)-(d) of Figure 8 depict the $\tau_{diff}$ in red next to the $\tau^{(1,2)}_{peaks}$ curves. The key difference is in the use of the thermal ensemble average which employs all the eigenvalues to compute the specific heat.
It is crucial to identify the $\tau$-scale that resolves the phase transition from the single trivial scale of the star graph to the more complex topological traits of the amorphous network depicted in Figure 5.(b). We consider such resolution scale of significant importance for the computation of the efficiency $\eta$ across the transition as it defines an "ultra-violet" cut-off in the characteristic frequency of the network across the transformation, similarly to what was discussed in the context of the Laplacian-Renormalization group [3]. In Section S3 of the Supplementary Material the efficiency, $\eta$, is computed in the range interval $\tau<10$ as is indicated by the diffusion time defined in [1] $\tau_{diff}=1/\lambda_{2}$, i.e. the red curves depicted in panels (a). Whether computed at those scales, $\eta$ shows a change in concavity and no relation with the pattern-matching accuracy.
Finally, note in Figure 8 the slow down of diffusion due to the increasing "volume" of the network discussed in Section 4 and indicated by the $\tau^{(2)}_{peak}$ curve as a function of $r$. We remark that $\tau^{(2)}_{peak}$ is a function of the parameter $r$ only, and the latter shapes the degree distribution only, as the number of nodes and edges is fixed across the transition. We note that the Portegy’s algorithm stays in the regime of sparsity, symbolically $\delta(\frac{N-1}{|E|}-1)$.

### v5

@@ II.2.1 Network growth process
Figure 2: Search method based on relative distance. In (a) the growth process is illustrated. At each step, a new pattern is linked to either the current parent pattern, $P$, or to one of its children, $C_{i}$. The parameter $r$ modulates the region of attraction of the child, indicated by the gray circle around $C_{1}$, and the one of the parent node. Points falling within the distance $r\cdot d(P,C_{i})$ are linked to the child, otherwise, an edge will be defined between the parent node, $P$, and the new pattern. In (b) the search method based on relative distance is illustrated. Once a search pattern arrives at node $P$ it must choose to search either the subtree of $C_{1}$ or $C_{2}$. The choice is based on minimizing the distance to the expected patterns contained in the subtree of $C_{i}$. In the depicted case, patterns closer to the search pattern, $S$, are expected to be in the subtree of $C_{2}$. Adapted from [58].
II.2.2 Search phase
@@
After the tree formation described in Sec. II.2.1, the Portegys algorithm enters the search phase, where the best match to an unknown search pattern, $S$, is given using a first-best approach. The original goal of this algorithm was to restructure a list of examples in a tree-based way, such that the matching accuracy would not be much worse than nearest-neighbor matching on a full list while providing a considerable speed up due to the hierarchical organization into a tree-based structure. The growth process, described in Sec. II.2.1 and depicted in Fig. 2(a), ensures that subtrees of a given node have on average decreasing distances for decreasing hierarchical levels.
This provides the possibility of using a search technique based on relative distances. The search method is depicted in Fig. 2(b) and is an iterative process that begins at the root-node pattern. Once a search pattern, $S$, arrives at a certain parent node, $P$, it must decide to continue to search for the best match into one of its children subtrees, $C^{\prime}_{1}$, $C^{\prime}_{2}$ in Fig. 2(b). The choice is taken minimizing the Portegys search-distance: $d(S,C_{i})_{search}=d(S,C_{1})-r\cdot d(P,C_{i})=d(S,C^{\prime}_{i})$, which quantifies the distance between the $S$-pattern and the potential patterns appearing in the subtree of each child, $C_{i}$, collectively indicated by the gray circle, $C^{\prime}_{i}$, in Fig. 2(b). Where $d(\cdot)$ indicates any metric distance equipped with semi-positivity, symmetry, and triangular inequality properties. If $d(S,C_{i})_{search}$ is less than zero then it is set to zero. In Fig. 2(b) the bold segments in red indicate the search distance to each child node and equivalently the expected distance between the unknown pattern $S$ and their expected neighborhoods of patterns $C^{\prime}_{i}$.
II.3 Datasets
Our results are evaluated over subsets of four different datasets ranked based on their complexity (Sec. III.1): the NIST [59], the MNIST [60], the FashionMNIST [61], and the CIFAR10 [62] datasets. These datasets contain randomly collected instances of image samples out of 10 classes. It should be noted that in this type of data, a pattern class can generally not be represented by a single centroid vector, but will contain multiple densities (“sub-style patterns”) such that the sample point cloud can be ordered naturally in a network structure.
The MNIST database contains $24\times 24$ grey-scale images of handwritten digits [0-9] commonly used to benchmark machine- and deep-learning algorithms. The NIST dataset is the ancestor of the MNIST, it contains $16\times 16$ black-and-white images of the handwritten digits [0-9]. The FashionMNIST is a large dataset of $28\times 28$ gray-scale images of fashion products from 10 categories. The CIFAR-10 dataset contains $32\times 32$ color images belonging to 10 classes. In all our experiments a subset of 2000 samples is extracted from the train set and the test set independently. Samples from the FashionMNIST and MNIST are always binarized. The samples from the CIFAR10 are normalized between 0 and 1 dividing each pixel value by 255, and values are rounded up to eighth decimal. In the Supplemental Material [63], we show results for larger networks generated from $4000$ samples from the MNIST and FashionMNIST. The 10 classes are equally represented in both the test and training set.
III Results
III.1 Dataset complexity
The datasets used in our experiments can be ranked based on their estimated complexity, see Fig. 3. We estimate each dataset’s complexity with multiple methods on the same subset of $2000$ training samples used in the rest of the manuscript.
The first heuristic approach is based on the test accuracy of a small multilayer perception (MLP) (2 layers with 128 units each). In pattern recognition, the complexity of a dataset is determined by the complexity of the curvilinear hyperplanes that would be needed to optimally separate patterns from different classes. A simple but effective parameter to gauge this facet of complexity is the accuracy that a classifier (such as the MLP) can achieve in determining the true class label for a data point. Thus, a more complex dataset would result in a lower accuracy for neural networks of fixed size when compared to a less complex dataset.
The second method we use is the computation of the Lempel-Ziv complexity [45] of the aggregated training set. In this case, the samples need to be binarized.
The third method concerns the measure of the intrinsic dimension [43, 48] of the dataset using the Maximum Likelihood Estimation approach introduced in Ref. [46] with the correction proposed in Ref. [47]. A given dataset of gray-scale images of size $pixel\times pixel$ a priori lies on a high-dimensional space of $\mathbb{R}^{pix\times pix}$, nevertheless, because it might contain a specific class of objects that account for the “world” of our problem, for instance, handwritten digits, it spans a lower dimensional manifold. Even though the definition of such a manifold is an open problem, it is possible to estimate its dimension based on the distance of the $n$-neighbouring points [43, 48]. We set the hyper-parameter $n=3$.
Figure 3: Dataset complexity is estimated with multiple methods for each training dataset of 2000 samples encountered in this work. In green an estimate is given computing the Lempel-Ziv complexity [45], while in orange the intrinsic dimension is estimated with the Maximum Likelihood Estimation approach [46, 47, 48]. In blue the accuracy on the test set for a multilayer-perceptron (MLP) with two hidden layers of 128 units each. The left axis refers to the quantities in green and orange, while the right axis refers to the multilayer-perceptron accuracy. The CIFAR10 dataset is estimated as the most complex, while the NIST dataset is the least complex among the four.
III.2 Pattern-matching efficiency
Mapping a high-dimensional vector collection into a hierarchical tree structure limits the number of comparisons needed to match an unknown pattern to a prototypical example. Namely, the algorithm introduced by Portegys [27, 28] is equipped with a pattern-searching method illustrated in Fig. 2 and described in Sec. II.2.2, representing the functional aspect of the system.
@@ III.2 Pattern-matching efficiency
We define a pattern-matching efficiency, $\theta$, in terms of class match accuracy, $A$, and the number of distances computed, $N_{dist}$:
@@
the logarithm is taken since the range of $N_{dist}$ can virtually span the interval $N_{dist}\in(0,N]$, where $N$ is the number of nodes. For the case of a star graph, $r\rightarrow 0$, the number of computed distances is maximum $N_{dist}=N$; the number of computed distances decreases for increasing $r$, potentially lowering the class-match accuracy. Nevertheless, we highlight Fig. 4(d) and note that the CIFAR10 dataset exhibits the global accuracy peak at an intermediate value of $r$. We interpret this observation by saying that for some datasets (see also the case of the FashionMNIST in Sec. S7 of the Supplemental Material [63]) specific network structures do possess the ability to better encode information of the “world” in their traits. We postpone the discussion of this finding to Secs. III and IV.
When defining the pattern-matching efficiency, we are making the simplified assumption that the pattern-matching algorithm runs on a physical substrate such that visiting any node has a fixed cost in terms of energy or in terms of general resources. Hence, the efficiency, $\theta$, quantifies the trade-off between the cost of used resources and the gain in performance, as depicted in Fig. 4(c),(f) for the MNIST and CIFAR10 respectively.
While in Sec. III.3 we discuss the full range of $r$. In the rest of the manuscript, we limit the discussion to $r\leq 0.85$ as the accuracy, and hence the pattern-matching efficiency, quickly drops for larger $r$, moreover, the algorithm shows instability as it doesn’t necessarily converge for $r\geq 0.9$ and $N=2000$.
@@
Figure 4: Pattern-matching efficiency in (c),(f) measured over the NIST and CIFAR10 respectively. The efficiency, $\theta$, is defined as the trade-off between the classification accuracy, (a),(d), and the logarithm of the number of distances computed to find the best match, (b),(e). The dashed-red vertical line indicates the optimal $r$ value in terms of efficiency, $\theta$. The shaded grey area designates the standard deviation. The purple shaded area in panel (c),(f) indicates a tolerance $r$-domain of optimal efficiency. The tolerance domain is defined as the $r$-values such that the corresponding $\theta$-values are included in one standard deviation computed at the maximum $\theta$, whose y-axis projection is indicated by the yellow shaded area. Curves are computed over 50 independent simulations with a subset of 2000 samples extracted from the NIST and CIFAR10 datasets. The 10-digit classes are equally represented both in the train- and test sets.
III.3 Phase portrait
@@
[…]
The Portegys algorithm produces three prototypical structures as those depicted in Fig. 1 and Fig. 5(b), depending on the value of the control parameter, $r$ (see in Methods Sec. II.2).
For $r\rightarrow 0$, every node is linked to the root node and a star graph turns out (see in Fig. 1 the central node, which is the first node selected in the growth process). For some intermediate value of $r$, a ragged tree structure with progressively smaller distances between the patterns is produced; while for $r\rightarrow 2$, a 1d-lattice is created.
[…]
@@
The second change of phase is the transition from the star graph to the amorphous configuration at $r\simeq 0.35$. Here, the emergence of scale-invariant structural properties is reflected by the progressively larger specific heat plateau in between the two $C$-peaks, as shown in Fig. 5(a),(c). For $r\rightarrow 2$, the plateau would include the full 1d-lattice in the thermodynamic limit $N\rightarrow\infty$.
@@
Before the transition, a single peak in the specific heat identifies a single trivial $\tau$-resolution scale below which the informationally connected components are the single nodes, and after which the informationally connected component is the full graph. After the phase transition, a second peak of the specific heat emerges at larger $\tau$ scales. This second peak of $C$ in Fig. 5(c) identifies the scale at which the information is integrated between the distinct modules. In other words, the second peak reflects the “degree” of the large-scale modules, in much analogy to the first peak, which reflects the presence of hubs at the local connectivity scale of the nodes. Note that we exclude both the clusters, i.e. triangles, and higher-order cycles from our discussion, as the tree networks we address are purely loop-less.
@@
The star graph and amorphous network phase space regions are those investigated in this work. Because the peaks in the thermodynamic and pattern-matching efficiencies always occur in the amorphous and scale-invariant region of the phase plane. In some cases, they appear in the vicinity of the transition border between the two at $r\simeq 0.35$. This transition, conversely to what we did with the amorphous to regular-lattice transition discussed above, can’t be identified by a quantity resolving the large-scale property of the network. Indeed, we observe in Fig. 6(b) the lack of a peak in the variance of $\hat{d}$, which would indicate no sign of bona fide phase transition. Nevertheless, a phase transition exists and an appropriate order parameter needs to be employed to identify it.
@@
This phase change occurs at smaller resolution scales with respect to the first phase transition we discussed. Specifically, it involves the mesoscales properties changing from the star graph to the scale-invariant regime (see the bifurcation of the specific heat in Fig. 5(a),(c)). Thus, we can identify this transition using as an order parameter the modularity [65], $Q$, of the tree which we show in Fig. 6(a). We are taking advantage of the fact that the modularity for trees is bounded between $[0,1]$ and that in the case of a 1d-lattice, the modularity is equal to $1$ if we partition the network into $N$ communities, which is the case if we use the community detection algorithm introduced in Ref. [66]. We stress that modularity is a mesoscale descriptor while the diameter of the network is a large-scale descriptor. Thus, a mesoscale resolution is required to monitor the structural changes occurring across the star graph phase to the amorphous scale-invariant phase transition. This motivates the choice of the mesoscale $\tau\sim 10$ employed in Sec. III.
@@
As a second possible order parameter, for the mesoscale phase-transition, we show in Fig. 6(c) the difference of entropy at the two $\tau^{(1,2)}_{peak}$ specific heat peaks. This order parameter counts the normalized (log)number of informationally connected nodes across the $\tau$-scales of the specific heat plateau. In simple terms, we count the fraction of nodes included in the scale-invariant region. We observe a discontinuity at the border of the scale-invariant transition that is not observed at the border of the amorphous to regular lattice transition.
@@
As already discussed in Sec. II.1.1, also along the $\tau$-direction an informational phase transition exists [34, 35] delimited by the peaks of $C$ at $\tau^{(1,2)}_{peak}$. Using the specific heat peaks, we can distinguish three dynamical phases (see Fig. 5(d)). For $\tau<\tau^{(1)}_{peak}$, the network is in the informationally segregated phase; while for $\tau>\tau^{(2)}_{peak}$ is in the informationally integrated phase. Only for $r>0.35$ and $\tau^{(1)}_{peak}<\tau<\tau^{(2)}_{peak}$, we observe the dynamical regime of constant-rate entropy loss characteristic of scale-invariant structures with dilation symmetry [35, 36].